a Father has two sons. The oldest son is 3 yrs older then the younger son. The father is 6 times older then the oldest son. sum of the ages of the 3 people is 47. Calculate the ages of the 3 individuals

Answer by ReadingBoosters(3246)

x = father
y = older son
z = younger son
y = 3 + z
x = 6(y)
x + y + z = 47
Plug the first two equations into the third equation
6y + (3 + z) + z = 47
6(3+z) + 3 + z + z = 47
18 + 6z + 3 + z + z = 47
21 + 8z = 47
21 - 21 + 8z = 47 - 21
8z = 26
8x/8 = 26/8
z = 3.25 youngest son
y = 3 + z = 3 + 3.25 = 6.25 oldest son
x = 6(6.25) = 37.5 father